# Arithmetical Complexity and Absoluteness of Rigidity Phenomena for Ulam Sequences

### v1

@@ Abstract
We analyse the logical complexity and absoluteness of a range of natural statements about Ulam sequences and Ulam sets, with particular emphasis on the rigidity phenomena introduced by Hinman, Kuca, Schlesinger and Sheydvasser. For each pair of coprime integers $a<b$ we view the associated Ulam sequence $U(a,b)$ as a recursive subset of $\mathbb{N}$ and consider expansions of the form $(\mathbb{N},+,\mathrm{U}_{a,b})$.
Our first main result is a uniform coding of Ulam sequences and of the “interval with periodic mask” patterns appearing in rigidity conjectures into first-order arithmetic. Using this, we show that the strongest natural rigidity statements, regularity statements (eventual periodicity of gaps), and density statements for $U(a,b)$ are all arithmetical and lie at low levels of the arithmetical hierarchy (e.g. $\Sigma^{0}_{2}$ or $\Pi^{0}_{3}$).
As a consequence, these statements are absolute between transitive models of $\mathrm{ZFC}$ with the same $\omega$, and in particular their truth value is insensitive to forcing extensions and to additional set-theoretic axioms such as $\mathrm{CH}$ or large cardinal hypotheses. We also initiate a model-theoretic study of the expansions $(\mathbb{N},+,\mathrm{U}_{a,b})$, relating combinatorial rigidity of $U(a,b)$ to potential tameness properties (such as the absence of the independence property) and to interpretability of multiplication.
@@ 1.1 Ulam sequences and rigidity
A second strand of work has highlighted striking global patterns in Ulam sequences and related objects. Steinerberger [Steinerberger2017HiddenSignal] observed that for a certain explicit real number $\alpha$, the sequence of fractional parts $\{u_{n}\alpha\}$, where $(u_{n})$ enumerates $U(1,2)$, exhibits a highly non-uniform limiting distribution concentrated in the middle third of $[0,1]$. This “hidden signal” is especially surprising in light of Weyl’s equidistribution theorem, which asserts that for any sequence of distinct integers $(a_{n})$ and Lebesgue-almost every $\alpha$, the sequence $\{a_{n}\alpha\}$ is equidistributed. At the same time, Kravitz and Steinerberger [KravitzSteinerberger2018] introduced higher-dimensional Ulam sets in $\mathbb{Z}^{n}$ and observed rigid geometric patterns in numerical experiments.
@@
Building on these observations, Hinman, Kuca, Schlesinger and Sheydvasser [HinmanKucaSchlesingerSheydvasser2019JNT, HinmanKucaSchlesingerSheydvasser2019Involve] developed a series of rigidity results for Ulam sequences and Ulam sets. Roughly speaking, they showed that as one varies the initial parameters (for instance, the second term $b$ in $U(1,b)$), large initial segments of the resulting sequences can be described by a finite amount of data that does not depend on $b$ itself, but only on its residue class modulo some fixed $L$. More concretely, for families $U(a,b)$ they produce decompositions of $U(a,b)\cap[1,Cb]$, valid for all sufficiently large $b$ in a fixed congruence class, as finite unions of “intervals with periodic masks” with endpoints depending linearly on $b$. Similar phenomena occur for higher-dimensional Ulam sets, where one sees only finitely many possible large-scale shapes up to a natural notion of equivalence. These results support the view that Ulam sequences, while seemingly irregular at small scales, exhibit a surprising degree of large-scale order.
1.2 A logical perspective
@@ 1.3 Main results
is equivalent to a $\Sigma^{0}_{2}$ sentence of first-order arithmetic. At the family level, the statement that every coprime pair $(a,b)$ satisfies strong rigidity,
@@
is shown in Theorem 4.4 to be equivalent to a $\Pi^{0}_{3}$ sentence.
@@
The detailed statements and proofs of these results are given in Sections 37 below.
@@ 1.4 Organisation of the paper
In Section 2 we recall the definition of Ulam sequences and Ulam sets, review basic facts about the arithmetical hierarchy and absoluteness, and fix a uniform coding of finite combinatorial data by natural numbers. Section 3 is devoted to the uniform arithmetisation of Ulam sequences and rigidity patterns: we construct the predicate $\mathrm{U}(a,b,m)$ and the pattern predicates $\mathrm{Pattern}(c)$ and $\mathrm{InPattern}(c,a,b,m)$, and prove Theorem 3.2. In Section 4 we analyse the logical complexity of strong rigidity statements, both for fixed $(a,b)$ and at the level of families, establishing the rigidity complexity results of Theorems 4.2 and 4.4. Section 5 treats regularity and density: we show that eventual periodicity of gaps, finiteness-of-evens, and various density assertions are all low in the arithmetical hierarchy, proving Theorems 5.3 and 5.8. In Section 6 we draw the resulting absoluteness and proof-theoretic consequences for these arithmetical statements. Section 7 initiates a model-theoretic study of the expansions $(\mathbb{N},+,\mathrm{U}_{a,b})$, proving Proposition 7.7 and formulating several conjectures relating combinatorial rigidity to tameness properties. We conclude in Section 8 with further directions and open problems.
@@ 2 Preliminaries
2.1 Ulam sequences and Ulam sets
@@
We now briefly recall the notion of Ulam sets in higher dimensions. Following Kravitz and Steinerberger [KravitzSteinerberger2018] and Hinman–Kuca–Schlesinger–Sheydvasser [HinmanKucaSchlesingerSheydvasser2019Involve], we work in $\mathbb{Z}^{n}$ equipped with a fixed norm $\|\cdot\|$ and the induced well-ordering of $\mathbb{Z}^{n}_{\geq 0}$ by nondecreasing norm, breaking ties lexicographically.
Definition 2.2.
Let $n,k\geq 1$, let $v_{1},\dots,v_{k}\in\mathbb{Z}^{n}_{\geq 0}$ be distinct initial vectors, and fix a norm $\|\cdot\|$ on $\mathbb{R}^{n}$. The $(k,n)$-Ulam set $U(v_{1},\dots,v_{k})\subseteq\mathbb{Z}^{n}_{\geq 0}$ is defined recursively by:
  • initially, $v_{1},\dots,v_{k}\in U(v_{1},\dots,v_{k})$;
  • at each step, among all $w\in\mathbb{Z}^{n}_{\geq 0}\setminus U(v_{1},\dots,v_{k})$ that can be written as $w=x+y$ in exactly one way with $x,y\in U(v_{1},\dots,v_{k})$ and $x\neq y$, we adjoin to $U(v_{1},\dots,v_{k})$ the element $w$ of minimal norm (breaking ties lexicographically).
When $n=1$ and $k=2$ with $v_{1}=a$, $v_{2}=b$, and the usual absolute value norm, this definition recovers $U(a,b)$ (up to a harmless choice of tie breaking). In dimension $n=2$, the resulting $(k,2)$-Ulam sets often exhibit striking geometric structure: for example, when $k=3$ and the initial vectors are $(1,0)$, $(0,1)$ and $(v_{1},v_{2})$ with $v_{1},v_{2}>0$, one sees only a small finite list of possible large-scale shapes, up to symmetry and finitely many exceptions [KravitzSteinerberger2018, HinmanKucaSchlesingerSheydvasser2019Involve].
In the present paper we will use Ulam sets primarily as motivation and as a source of examples. Our main focus will be on the one-dimensional sequences $U(a,b)$ and on viewing them as definable subsets of $\mathbb{N}$ inside the structures $(\mathbb{N},+)$ and $(\mathbb{N},+,\mathrm{U}_{a,b})$.
@@ 2.2 The arithmetical hierarchy and absoluteness
This is standard. The satisfaction relation for arithmetical formulas is defined by primitive recursion on the complexity of formulas, using only quantification over $\mathbb{N}$. Since $M$ and $N$ have the same underlying structure (,+,×,)($\mathbb{N}$$\times,\leq$), they compute the same truth values for all arithmetical formulas. In particular, they agree on all arithmetical sentences. ∎
@@ 2.3 Coding finite data by natural numbers
Remark 2.7.
For the specific family $U(1,n)$, the pattern components considered above specialise to the “piecewise linear” descriptions that have been computed experimentally in the literature. In particular, Sheydvasser has implemented algorithms which, given the numerical data of $U(1,2),\dots,U(1,14)$, recover coefficients $a_{i},b_{i},c_{i},d_{i}$ such that for large $n$ one has a decomposition of the form
$U(1,n)\cap[1,Cn]\;=\;\bigcup_{i<k}\bigl([a_{i}n+b_{i},\,c_{i}n+d_{i}]\cap\mathbb{N}\bigr),$
possibly together with periodic masks inside each interval; see [Sheydvasser2021LinearPolynomials, SheydvasserUlamCode]. In our notation, these intervals correspond to components with $A_{1}=0$, $A_{2}=a_{i}$, $B_{1}=0$, $B_{2}=c_{i}$ and suitable constants $p=b_{i}$, $q=d_{i}$, together with the trivial mask $S=\{0,\dots,L-1\}$ (or a non-trivial mask when internal periodicity is present). Thus the notion of pattern code used here encompasses the concrete forms of the decompositions that arise in computational work on Ulam sequences.
@@ 3.2 Coding patterns for rigidity
In our notation, this single block is captured by a pattern component of the form
$\bigl([A(1,n)n+p,\,B(1,n)n+q]\cap\mathbb{N},\;L,\;S\bigr),$
with
$A(1,n)=4,\quad B(1,n)=5,\quad p=2,\quad q=-1,\quad L=1,\quad S=\{0\}.$
The associated contribution to $\mathrm{PatternSet}_{c}(1,n)$ is then
$\{m\in\mathbb{N}:4n+2\leq m\leq 5n-1\},$
@@ 5.3 Density
Strong rigidity $\mathrm{Rig}(a,b)$ $\Sigma^{0}_{2}$ (Thm. 4.2) $\Pi^{0}_{3}$ (Thm. 4.4)

### v2

@@ Abstract
We analyse the logical complexity and absoluteness of natural statements about Ulam sequences, with particular emphasis on the rigidity phenomena introduced by Hinman, Kuca, Schlesinger and Sheydvasser for the family $U(1,n)$. For each pair of coprime integers $a<b$ we view the associated Ulam sequence $U(a,b)$ as a recursive subset of $\mathbb{N}$ and consider expansions of the form $(\mathbb{N},+,\mathrm{U}_{a,b})$.
Our first main result is a uniform coding of Ulam sequences and of the interval-with-periodic-mask patterns appearing in rigidity conjectures into first-order arithmetic. Using this, we show that strong rigidity, regularity (eventual periodicity of gaps), and density statements for $U(a,b)$ are all arithmetical and lie at low levels of the arithmetical hierarchy (specifically $\Sigma^{0}_{2}$ or $\Pi^{0}_{3}$).
As a consequence, these statements are absolute between transitive models of $\mathrm{ZFC}$ with the same natural numbers: their truth value cannot be changed by forcing, and is independent of the Continuum Hypothesis and large cardinal axioms. We also study the expansions $(\mathbb{N},+,\mathrm{U}_{a,b})$ model-theoretically, showing that combinatorial rigidity implies tameness (NIP, dp-minimality, non-interpretability of multiplication).
@@ 1.1 Ulam sequences and rigidity
A second strand of work has highlighted striking global patterns in Ulam sequences. Steinerberger [Steinerberger2017HiddenSignal] observed that for a certain explicit real number $\alpha$, the sequence of fractional parts $\{u_{n}\alpha\}$, where $(u_{n})$ enumerates $U(1,2)$, exhibits a highly non-uniform limiting distribution concentrated in the middle third of $[0,1]$. This “hidden signal” is especially surprising in light of Weyl’s equidistribution theorem, which asserts that for any sequence of distinct integers $(a_{n})$ and Lebesgue-almost every $\alpha$, the sequence $\{a_{n}\alpha\}$ is equidistributed.
@@
Building on these observations, Hinman, Kuca, Schlesinger and Sheydvasser [HinmanKucaSchlesingerSheydvasser2019JNT, HinmanKucaSchlesingerSheydvasser2019Involve] developed a series of rigidity results for Ulam sequences. Roughly speaking, they showed that as one varies the second term $n$ in $U(1,n)$, large initial segments of the resulting sequences can be described by a finite amount of data that does not depend on $n$ itself, but only on its residue class modulo some fixed $L$. More concretely, for the family $U(1,n)$ they produce decompositions of $U(1,n)\cap[1,c_{j}n+d_{j}]$ (where $c_{j}\to\infty$), valid for all sufficiently large $n$ in a fixed congruence class, as finite unions of intervals with endpoints depending linearly on $n$.
Remark 1.1 (Status of rigidity for $U(a,b)$ with $a>1$).
The rigidity results of [HinmanKucaSchlesingerSheydvasser2019JNT] are established for the family $U(1,n)$. For sequences $U(a,b)$ with $a>1$, computational evidence collected since the original rigidity paper suggests that strong rigidity in the same form does not hold in general, although the correct conjectural description remains unclear (A. Sheydvasser, personal communication, 2025). We therefore focus primarily on the $U(1,n)$ family when discussing proven rigidity phenomena, while formulating our arithmetical complexity results for general coprime pairs $(a,b)$. The complexity classifications themselves (Theorems 4.35.8) hold regardless of whether the corresponding rigidity or regularity statements are true.
These results support the view that Ulam sequences, while seemingly irregular at small scales, exhibit a surprising degree of large-scale order—at least in the $U(1,n)$ family.
1.2 Main results and significance
The central contribution of this paper is to determine the logical complexity of statements about Ulam sequences and to draw consequences from this classification. Our main theorem, stated informally, is:
Main Theorem (Informal). The strong rigidity conjecture, gap regularity, and density statements for Ulam sequences $U(a,b)$ are all arithmetical statements of low complexity (specifically $\Sigma^{0}_{2}$ or $\Pi^{0}_{3}$). As a consequence:
  1. 1.
    Adding or removing set-theoretic axioms such as the Axiom of Choice, the Continuum Hypothesis, or large cardinal axioms does not change their truth values.
  2. 2.
    No set-theoretic forcing extension of $\mathrm{ZFC}$ can make one of these statements true while another such extension (with the same natural numbers) makes it false.
  3. 3.
    If $U(a,b)$ satisfies gap regularity, then the expansion $(\mathbb{N},+,\mathrm{U}_{a,b})$ is model-theoretically tame: NIP, dp-minimal, and does not interpret multiplication.
Why this matters: for number theorists
The rigidity conjectures for Ulam sequences are purely combinatorial questions about sequences of integers. It is natural to ask whether sophisticated set-theoretic machinery might help resolve them—perhaps the answer depends on whether we accept certain axioms, or perhaps forcing techniques could shed light on the problem.
Our results show definitively that this cannot happen. The statements in question are absolute: their truth or falsity is already determined by the structure of the natural numbers, and no set-theoretic axiom or technique can change the answer. This is both reassuring (the answer does not depend on one’s foundational commitments) and sobering (there are no set-theoretic shortcuts to resolving these combinatorial questions).
More specifically, we show that the key conjectures can be expressed using only two or three alternations of quantifiers over natural numbers. This places them at a very low level of logical complexity—comparable to statements like “there exist infinitely many twin primes” or “Goldbach’s conjecture holds.” The classification gives a precise sense in which these problems, while difficult, are “elementary” from a logical standpoint.
Why this matters: for model theorists
Ulam sequences provide a natural family of recursively defined predicates on $\mathbb{N}$ whose model-theoretic behavior is controlled by their combinatorial structure. The expansion $(\mathbb{N},+,\mathrm{U}_{a,b})$ sits between two well-understood extremes: Presburger arithmetic $(\mathbb{N},+)$, which is decidable and tame, and full arithmetic $(\mathbb{N},+,\times)$, which is undecidable and wild.
Under the assumption of combinatorial rigidity, we show that these expansions remain on the tame side: they are definitional expansions of Presburger arithmetic, hence NIP and dp-minimal. Without assuming rigidity, the situation is open and provides concrete test cases for questions about expansions of $(\mathbb{N},+)$ by sparse additive sets. The paper initiates a systematic study of these expansions and formulates precise conjectures about their tameness.
1.3 Logical background for non-logicians
For readers unfamiliar with mathematical logic, we provide informal explanations of the key concepts that appear in our results. Precise definitions are given in Section 2.
The arithmetical hierarchy
The arithmetical hierarchy classifies statements about natural numbers by the pattern of quantifiers (“for all” and “there exists”) needed to express them. The notation $\Sigma^{0}_{n}$ and $\Pi^{0}_{n}$ indicates $n$ alternating blocks of quantifiers:
$\Sigma^{0}_{1}$: $\exists x\,\varphi(x)$ where $\varphi$ is “directly checkable”
$\Pi^{0}_{1}$: $\forall x\,\varphi(x)$ where $\varphi$ is “directly checkable”
$\Sigma^{0}_{2}$: $\exists x\,\forall y\,\varphi(x,y)$
$\Pi^{0}_{2}$: $\forall x\,\exists y\,\varphi(x,y)$
$\Sigma^{0}_{3}$: $\exists x\,\forall y\,\exists z\,\varphi(x,y,z)$
Lower levels correspond to “simpler” statements. For example, “$U(1,2)$ has infinitely many even elements” is $\Pi^{0}_{2}$ (for all $N$, there exists an even $m>N$ in $U(1,2)$), while “the gap sequence of $U(a,b)$ is eventually periodic” is $\Sigma^{0}_{2}$ (there exist $N$ and $p$ such that for all $k\geq N$, the gaps repeat with period $p$).
Absoluteness
A statement is absolute if its truth value does not depend on the ambient set-theoretic universe. Different models of set theory (ZFC) can disagree about many things—the size of the continuum, whether certain large cardinals exist, etc.—but they all agree on arithmetical statements, because they all contain the same natural numbers with the same addition and multiplication.
The practical consequence: if you prove that a conjecture about Ulam sequences is arithmetical (as we do), then you know immediately that the Axiom of Choice, the Continuum Hypothesis, large cardinal axioms, and set-theoretic forcing cannot affect its truth value.
NIP and tameness
In model theory, structures are classified by how “wild” or “tame” their definable sets are. A theory is NIP (“not the independence property”) if its definable families of sets cannot encode arbitrary combinatorial patterns—roughly, there is a uniform bound on how complicated the interactions between definable sets can be.
Presburger arithmetic $(\mathbb{N},+)$ is a paradigm of tameness: NIP, decidable, and with a complete understanding of definable sets (they are “semilinear”). Full arithmetic $(\mathbb{N},+,\times)$ is wild: it can encode arbitrary computation, and its definable sets are extremely complicated.
When we say that $(\mathbb{N},+,\mathrm{U}_{a,b})$ is “tame” under rigidity assumptions, we mean it inherits the good behavior of Presburger arithmetic rather than the pathologies of full arithmetic.
1.4 Technical overview
@@ 1.5 Main results
is equivalent to a $\Sigma^{0}_{2}$ sentence of first-order arithmetic. At the family level, the conjectural statement that every coprime pair $(a,b)$ satisfies strong rigidity,
@@
is shown in Theorem 4.6 to be equivalent to a $\Pi^{0}_{3}$ sentence. (As noted in Remark 1.1, this family-level conjecture is not expected to hold as stated; strong rigidity appears to fail for $U(a,b)$ with $a>1$. However, the complexity classification remains valid regardless of the statement’s truth value.)
@@
The detailed statements and proofs of these results are given in Sections 37 below. We emphasise that, beyond the $U(1,n)$ family where rigidity results are known, our theorems concern the logical complexity and absoluteness of the corresponding rigidity, regularity, and density schemas for general coprime $(a,b)$; we do not claim these combinatorial statements hold in full generality.
@@ 1.6 Organisation of the paper
Section 1.3 together with the technical overview provide informal and formal background for the main results. Section 2 gives precise definitions: Ulam sequences, the arithmetical hierarchy, absoluteness, and the coding of finite combinatorial patterns. Section 3 constructs the uniform $\Delta^{0}_{1}$ predicate $\mathrm{U}(a,b,m)$ and the pattern predicates used in rigidity statements (Theorem 3.2). Section 4 analyses the complexity of strong rigidity, establishing Theorems 4.3 and 4.6. Section 5 treats regularity and density, proving Theorems 5.3 and 5.8. Section 6 draws absoluteness and proof-theoretic consequences. Section 7 develops the model-theoretic analysis of $(\mathbb{N},+,\mathrm{U}_{a,b})$, proving Proposition 7.11 and formulating tameness conjectures. We conclude in Section 8 with open problems.
@@ 2 Preliminaries
2.1 Ulam sequences
@@
Remark 2.2 (Higher-dimensional generalisations).
Ulam-type constructions can be defined in higher dimensions: Kravitz–Steinerberger [KravitzSteinerberger2018] introduced Ulam sets in $\mathbb{Z}^{n}$, and Hinman–Kuca–Schlesinger–Sheydvasser [HinmanKucaSchlesingerSheydvasser2019Involve] established rigidity phenomena for these as well. The arithmetisation methods of this paper extend naturally to such settings, but we focus exclusively on the one-dimensional sequences $U(a,b)$ for concreteness.
@@ 2.2 The arithmetical hierarchy and absoluteness
The satisfaction relation for arithmetical formulas can be defined by primitive recursion on the complexity of formulas, using only quantification over $\mathbb{N}$. In particular, the truth value of an arithmetical sentence depends only on the underlying structure (,+,×,)($\mathbb{N}$$\times,\leq$). Since $M$ and $N$ have the same natural numbers and the same interpretations of $\times$$\leq,$ they assign the same truth value to every arithmetical sentence. ∎
@@ 2.3 Coding finite data by natural numbers
Remark 2.7 (The family $U(1,n)$: mask-free decompositions).
For the specific family $U(1,n)$, the pattern components considered above specialise to the “piecewise linear” descriptions that have been computed experimentally in the literature. In particular, Sheydvasser has implemented algorithms which, given the numerical data of $U(1,2),\dots,U(1,14)$, recover coefficients $c_{j},d_{j}$ (with $c_{j}\to\infty$ as $j\to\infty$) such that for all sufficiently large $n$ in a fixed residue class modulo some period $L$, one has a decomposition of the form
$U(1,n)\cap[1,c_{j}n+d_{j}]\;=\;\bigcup_{i<k}\bigl([a_{i}n+b_{i},\,a^{\prime}_{i}n+b^{\prime}_{i}]\cap\mathbb{N}\bigr);$
see [Sheydvasser2021LinearPolynomials, SheydvasserUlamCode]. Crucially, for the family $U(1,n)$ there are no masks: every interval $a_{i}n+b_{i},\,a^{\prime}_{i}n+b^{\prime}_{i}]$ contributes all of its integer points to $U(1,n)$, without periodic deletions. In our notation, this corresponds to pattern components with $L=1$ and $S=\{0\}$ (the trivial mask).
This mask-free property is specific to the $U(1,n)$ family. For general $U(a,b)$ with $a>1$, rigidity phenomena are less well understood, and non-trivial masks may in principle appear; see Remark 1.1 below. The general pattern-coding framework of Section 2.3 accommodates both cases.
@@ 3.2 Coding patterns for rigidity
In our notation, this single block can be realised by a single pattern component with coefficients
$A_{1}=0,\quad A_{2}=4,\quad B_{1}=0,\quad B_{2}=5,\quad p=2,\quad q=-1,\quad L=1,\quad S=\{0\}.$
For parameters $(a,b)=(1,n)$ this gives
$A(1,n)=A_{1}\cdot 1+A_{2}\cdot n+p=4n+2,\qquad B(1,n)=B_{1}\cdot 1+B_{2}\cdot n+q=5n-1,$
so the associated contribution to $\mathrm{PatternSet}_{c}(1,n)$ is exactly
$\{m\in\mathbb{N}:4n+2\leq m\leq 5n-1\}.$
@@ 4.1 Rigidity for a fixed $(a,b)$
Remark 4.2 (Initial-segment formulation).
For the family $U(1,n)$, rigidity results are often stated in an equivalent “initial-segment” form: there exist coefficients $c_{j},d_{j}$ (with $c_{j}\to\infty$ as $j\to\infty$) such that for all sufficiently large $n$ in a fixed residue class,
$U(1,n)\cap[1,c_{j}n+d_{j}]\;=\;\mathrm{PatternSet}_{c}(1,n)\cap[1,c_{j}n+d_{j}].$
This formulation better reflects how rigidity is computationally verified: one checks that $U(1,n)$ matches the predicted pattern on successively larger initial segments $1,c_{j}n+d_{j}]$. The two formulations are equivalent for our purposes, since the pattern set is eventually equal to the Ulam sequence if and only if they agree on all sufficiently large initial segments.
@@ 4.2 Family-level rigidity
Remark 4.5.
As discussed in Remark 1.1, the statement $\mathrm{FRig}$ is not expected to be true: computational evidence suggests that $\mathrm{Rig}(a,b)$ fails for many pairs with $a>1$. We analyse the complexity of $\mathrm{FRig}$ nonetheless, both because it illustrates the general method and because understanding why $\mathrm{FRig}$ is $\Pi^{0}_{3}$ clarifies the logical structure of family-level rigidity statements. A more refined family-level statement, restricting to $a=1$, would have the same complexity.
@@ 5.3 Density
Strong rigidity $\mathrm{Rig}(a,b)$ $\Sigma^{0}_{2}$ (Thm. 4.3) $\Pi^{0}_{3}$ (Thm. 4.6)